# Algorithmic Behaviors Across Regions: A Geolocation Audit of YouTube Search for COVID-19 Misinformation between the United States and South Africa

### v1

@@ 5 Labeling YouTube Videos
We obtained annotations for 3,075 videos, which were annotated by the first author and Amazon Mechanical Turk (AMT). The first author, as the expert, annotated 1,087 videos [^11] To gather more annotation, the first author followed Tomlein et al. (2021) and reannotated Serrano et al. (2020)’s YouTube video dataset, resulting in 143 more annotations.. For 1,988 videos, we obtained three AMT worker annotations per video. To assess the AMT workers’ agreement, we calculated the Fleiss’ Kappa Score and obtained $\kappa=0.62$, indicating “substantial agreement.” We employed the majority response to assign the final label, arriving at a final label for 1,899 videos. For the remaining 89 videos, all three AMT worker responses diverged. The first author annotated the 89 videos to obtain the final annotation values. In §11, we address the steps to minimize potential harm associated with exposing AMT workers to misinformation. See Appendix F for the AMT worker training, screening, compensation, and annotation task. Overall, the ground-truth dataset consisted of 3,075 videos, 820 of which were supporting, 837 opposing, 431 neutral, 409 irrelevant, 317 non-English, 228 on COVID-19 origins, and 33 URLs not accessible. To train our classifier, we excluded videos annotated as non-English and URL not accessible.
@@
To find the best-performing classifier, we systematically trained and evaluated various machine learning models, deep learning models, and LLMs on the three-class classification task, using combinations of text-based video metadata features such as titles, descriptions, transcripts, tags, and comments. Models included Support Vector Machine (Cortes 1995), XGBoost (Chen et al. 2016), DeBerta-v3-base and DeBerta-v3-large (He et al. 2021), and GPT-4 Turbo (v1106) (OpenAI et al. 2024). We held out 10% of the ground-truth dataset as our test set and used accuracy, weighted F1-score, and macro F1-score to evaluate the model performances. The remaining 90% of the dataset was employed for training and validation. See Appendix G for the input feature descriptions and training procedure.
@@
Our best-performing model on the three-class classification task achieves a comparable or even exceeds the performance of other models from prior studies (Ng et al. 2023; Hou et al. 2020; Serrano et al. 2020; Liaw et al. 2023). For example, Papadamou et al. (2022) developed a binary classifier that detected pseudoscientific videos with an accuracy of 0.79 and an F1-score of 0.74. Nevertheless, we acknowledge that this is not a perfect performance, reflecting the complex nature of identifying COVID-19 misinformation within videos and signaling the need for further research efforts. With a 15% error rate in our classifier, we took additional steps to validate the reliability of our findings due to potential labeling errors (see subsection “Validation of Results” in Appendix J).
@@ 6 Quantifying Misinformation Bias
To quantify the misinformation present in YouTube SERPs, we adopted the misinformation bias score metric developed by Hussein et al. (2020). The score determines the misinformation bias in a ranked list, giving more weight to the annotation labels of higher-ranked videos in the calculation: $\frac{\sum^{n}_{r=1}(x_{r}*(n-r+1))}{\frac{n*(n+1)}{2}}$, where $x$ represents the annotation label of the video, $r$ represents the ranking of the video, and $n$ represents the total number of videos in the SERP. To conform to the video annotation scale in Hussein et al. (2020), we mapped our 3-class labels from §5 to a normalized scale of -1, 0, and 1 based on their stance towards COVID-19 misinformation. Videos that oppose COVID-19 misinformation were assigned scores of -1, while those supporting it received a score of 1. However, videos that fell into the merged category, including irrelevant, neutral, and COVID-19 origins labels, do not support nor oppose COVID-19 misinformation. Thus, they [^13] We also assigned a 0 score to removed videos whose URL was inaccessible ($<$1% of data), providing a conservative estimate of the misinformation bias in SERPs. received a score of 0. Therefore, the misinformation bias score of a SERP is a continuous value ranging from -1 (all videos oppose misinformation) to +1 (all videos support misinformation). A positive score indicates a lean toward misinformation, while a negative score indicates a lean toward content opposing misinformation. A score of 0 suggests a set of content that neither supports nor opposes misinformation. A higher score suggests a higher prevalence of misinformation in the SERP.
@@ 7 Results
Here, we compare the prevalence of misinformation in YouTube SERPs across geolocations, topics, and filters. A test of normality revealed that our data is not normal. Thus, we used the non-parametric Mann-Whitney U Test for pairwise comparisons and Kruskal-Wallis tests, followed by the post-hoc Conover-Iman tests with Bonferroni adjustment for multiple comparisons. We provide comprehensive details of all significance tests conducted in Appendix I. Additionally, we include additional analyses on the misinformation bias scores in search queries and temporal trends of misinformation bias between the US and SA in Appendix J.
@@
Among the top-N videos, the top-10 videos in the SERPs have the highest misinformation bias scores. We observed that the top-10 videos in the SERPs have the highest misinformation bias scores for both the US and SA (see Figure 2). This suggests that misinformative content is more prevalent in top-10 search results than lower-ranked results. Within the top-10 search results, we observed that nearly a third of the contents were misinformative: 31.55% of the search results supported COVID-19 misinformation, 36.03% were opposing, and 32.42% belonged to the remaining classes.
Bots in SA encountered significantly more misinformative SERPs than bots in the US for both the top-10 and top-20 search results. For the top-10 videos, the effect size r was 0.49 (see Appendix Table 9), indicating a “medium effect” (Cohen 1988). This effect size indicates that our observed differences moderately carry practical significance, suggesting that geolocation influences the algorithmic behaviors of YouTube search — in our case, resulting in statistically significantly more misinformative SERPs in SA compared to the US. We discuss the potential implications of the observed differences in §8.
When considering lower-ranked search results, bots in the US encountered significantly more misinformative SERPs than bots in SA. This suggests that when considering the top-40 and top-50 videos in SERPs, bots in the US overall obtained more misinformative SERPs than bots in SA. However, it is important to note that 95% of user traffic is directed towards the first page of search results (Kaye 2013), where the top-10 results reside, making it more likely for users to interact with these results than with lower-ranked results. As a result, the difference in the top-40 and top-50 results may have less impact on user exposure compared to the top-10 results, where SA users are more likely to encounter misinformation. Since video viewers are most likely to engage with the top-10 videos in SERPs, our subsequent analysis focuses on these search results.
@@
Bots in the SA encountered significantly more misinformative SERPs than bots in the US for 3 topics. In Figure 3, the topics “5G claims” (r=0.80), “Bill Gates Claims” (r=0.44), and “Vaccine Content Claims” (r=0.59) exhibited statistically significant differences in misinformation bias between the US and SA, indicating medium to large effect sizes. These effect sizes imply practical significance, suggesting that users in SA may encounter significantly more misinformative SERPs than users in the US for these topics. Of particular concern is the presence of these differences in topics related to public health, such as “Vaccine Content Claims.“ In contrast, bots in the US encountered significantly more misinformative SERPs than bots in the SA for a single topic—“Spread of Virus.” We discuss the implications of these findings in §8.
@@
Topics “Biological Weapon” and “Lab Leak Theory” exhibited positive misinformation bias scores across all search filters in both countries (see Figure 4). Given that positive scores indicate that the SERPs lean toward misinformative content, this finding suggests the problematic nature of these topics, with bots exposed to misinformative SERPs regardless of their geolocations or the search filters used.
Topics “5G Claims,” “Spread of Virus,” and “Vaccine Content Claims” displayed negative bias scores across all search filters in both countries. This suggests that SERPs for these topics generally oppose misinformation in both the US and SA. This result may be attributed to YouTube’s targeted content moderation efforts on these topics (Hern 2020), potentially highlighting some successes in combating misinformation on its platform.
Three topics showed negative bias scores in SERPs sorted by “Relevance” and “View Count,” but positive scores when sorted by “Upload Date” and “Rating.” We observe this pattern for “Bill Gates Claims, “Treatment of Virus,” and “Population Control” topics, indicating that bots encounter more misinformative SERPs when searching for newly uploaded or highly rated videos for such topics.
@@
In contrast, “Relevance” and “View Count” filters showed negative bias scores in both countries. This suggests that SERPs sorted by relevant or most-viewed videos lean towards opposing misinformation. This may suggest that YouTube’s content moderation efforts prioritize relevant and highly viewed videos, which are likely to be surfaced by the search engine or have gained viewer attention.
Bots in SA received significantly more misinformative SERPs than bots in the US when sorting the result by “Relevance.” As shown in Figure 5, the effect size of r=0.86 indicates a large effect, suggesting that users in SA may encounter more misinformative SERPs than their US counterparts when using the default “Relevance” filter.
In contrast, bots in the US received significantly more misinformative SERPs than those in SA when results were sorted by “Rating.” The effect size r=0.48 indicates a medium effect. While this effect size is not as pronounced as when sorting by “Relevance” (r=0.86), it suggests that the observed differences have moderate practical significance. It is important to note that YouTube’s default filter is “Relevance,” suggesting that users are more likely to engage with SERPs sorted by this criterion. Consequently, the observed differences with the “Rating” filter may have a lesser impact on users compared to the “Relevance” filter, where bots in SA received significantly more misinformative SERPs than bots in the US.
@@ 8 Discussion
Auditing YouTube’s search engine for COVID-19 misinformation is urgently needed to ensure algorithmic accountability, protect public health, and develop more responsible web. Despite being designated a global pandemic in 2020, COVID-19 remains a public health threat (Colarossi 2024). While YouTube has been a valuable tool for finding health information during the pandemic, it has also disseminated harmful misinformation, impeding public health efforts and vaccine hesitancy (WHO 2022).
@@
Furthermore, our results indicate that bots in SA were exposed to significantly more misinformative SERPs compared to bots in the US for the topics “5G Claims,” “Bill Gates Claims,” and “Vaccine Content Claims” (Figure 3). Notably, South Africa experienced distressing incidents during the pandemic, such as the burning of 5G towers fueled by a conspiracy theory linking them to the spread of COVID-19 (Staff 2021). Concurrently, South Africa grappled with numerous public health challenges, including a vaccine supply shortage and a deep distrust of medical authorities contributing to vaccine hesitancy (Chutel and Fisher 2021). Given these circumstances, the heightened exposure of SA users to misinformation compared to US users on these topics could potentially exacerbate societal issues in SA, including vaccine hesitancy and undermining public health efforts.
Within the US, our analysis revealed that bots in Houston, Texas, received significantly more misinformative SERPs than those in Los Angeles, California. These observed geolocation differences within the US have important societal implications. A research study has highlighted that conservatives were more susceptible to believing falsehoods and conspiracy beliefs (Grabmeier 2021). Therefore, since Houston is in a conservative-leaning state (Texas), promoting significantly more misinformative search results to conservative-leaning users may potentially exacerbate the harm by reinforcing their beliefs in COVID-19 misinformation or leading them to adopt such beliefs.
@@
Historically, past audits have successfully generated awareness about these issues, creating pressure on platforms to improve their algorithm (Buolamwini 2019). Similarly, we hope our work drives meaningful changes on YouTube and inspires future audit research in the Global South.
@@ 9 Limitations and Future Directions
Our work is not without limitations. Our audit used only English search queries to maintain consistency in search query language across both countries. Future studies can audit YouTube’s search results by utilizing queries in diverse regional and local languages within the countries. Our study focuses on YouTube’s search engine; future studies can investigate the impacts of geolocation on the recommendation systems on the home page and video page.
Given the challenges of conducting audits in the Global South, our audit focuses on a country from the Global North and one from the Global South. Future works should conduct audits in multiple countries within the Global South. Our work uses the misinformation bias score (Hussein et al. 2020), capturing the amount of misinformation and the video’s rank but not its relevance. Future works can use metrics like the Convex Aggregated Measure (Lioma et al. 2017), which uses relevance and credibility in ranked lists.
Although we showed in §5 that our classifier performs comparably to YouTube models tackling related issues, we acknowledge that our classifier has an error rate of 15%, potentially influencing downstream misinformation analysis. We partly address this limitation by conducting additional analyses, which provided further evidence of the reliability of our findings despite the error rate of our classifier (see subsection “Validation of Results” in Appendix J).

### v2

@@ 5 Labeling YouTube Videos
We obtained annotations for 3,075 videos, which were annotated by the first author and Amazon Mechanical Turk (AMT). The first author, as the expert, annotated 1,087 videos. [^12] To gather more annotation, the first author followed Tomlein et al. (2021) and reannotated Medina Serrano, Papakyriakopoulos, and Hegelich (2020)’s YouTube video dataset, resulting in 143 more annotations. For 1,988 videos, we obtained three AMT worker annotations per video. To assess the AMT workers’ agreement, we calculated the Fleiss’ Kappa Score and obtained $\kappa=0.62$, indicating “substantial agreement.” We employed the majority response to assign the final label, arriving at a final label for 1,899 videos. For the remaining 89 videos, all three AMT worker responses diverged. The first author annotated the 89 videos to obtain the final annotation values. In §Ethical Statement, we address the steps to minimize potential harm associated with exposing AMT workers to misinformation. See Appendix F for the AMT worker training, screening, compensation, and annotation task. Overall, the ground-truth dataset consisted of 3,075 videos, 820 of which were supporting, 837 opposing, 431 neutral, 409 irrelevant, 317 non-English, 228 on COVID-19 origins, and 33 URLs not accessible. To train our classifier, we excluded videos annotated as non-English and URL not accessible.
@@
To find the best-performing classifier, we systematically trained and evaluated various machine learning models, deep learning models, and LLMs on the three-class classification task, using combinations of text-based video metadata features such as titles, descriptions, transcripts, tags, and comments. Models included Support Vector Machine (Cortes 1995), XGBoost (Chen and Guestrin 2016), DeBerta-v3-base and DeBerta-v3-large (He, Gao, and Chen 2021), and GPT-4 Turbo (v1106) (OpenAI et al. 2024). We held out 10% of the ground-truth dataset as our test set and used accuracy, weighted F1-score, and macro F1-score to evaluate the model performances. The remaining 90% of the dataset was employed for training and validation. See Appendix G for the input feature descriptions and training procedure.
@@
Our best-performing model on the three-class classification task achieves a comparable or even exceeds the performance of other models from prior studies (Ng, Hoffmann Pham, and Luengo-Oroz 2023; Hou et al. 2020; Medina Serrano, Papakyriakopoulos, and Hegelich 2020; Yi Liaw et al. 2023). For example, Papadamou et al. (2022) developed a binary classifier that detected pseudoscientific videos with an accuracy of 0.79 and an F1-score of 0.74. Nevertheless, we acknowledge that this is not a perfect performance, reflecting the complex nature of identifying COVID-19 misinformation within videos and signaling the need for further research efforts. With a 15% error rate in our classifier, we took additional steps to validate the reliability of our findings due to potential labeling errors (see subsection “Validation of Results” in Appendix K).
@@ 6 Quantifying Misinformation Bias
To quantify the misinformation present in YouTube SERPs, we adopted the misinformation bias score metric from Hussein, Juneja, and Mitra (2020). The score determines the misinformation bias in a ranked list, giving more weight to the annotation labels of higher-ranked videos in the calculation: $\frac{\sum^{n}_{r=1}(x_{r}*(n-r+1))}{\frac{n*(n+1)}{2}}$, where $x$ represents the annotation label of the video, $r$ represents the rank of the video in the SERP (e.g., $r=1$ indicates the top-most video in the SERP), and $n$ represents the total number of videos in the SERP. To conform to the video annotation scale in Hussein, Juneja, and Mitra (2020), we mapped our 3-class labels from §5 to a normalized scale of -1, 0, and 1 based on their stance towards COVID-19 misinformation. Videos that oppose COVID-19 misinformation were assigned scores of -1, while those supporting it received a score of 1. However, videos that fell into the merged category, including irrelevant, neutral, and COVID-19 origins labels, do not support nor oppose COVID-19 misinformation. Thus, they [^14] We also assigned a 0 score to removed videos whose URL was inaccessible ($<$1% of data), providing a conservative estimate of the misinformation bias in SERPs. received a score of 0. Therefore, the misinformation bias score of a SERP is a continuous value ranging from -1 (all videos oppose misinformation) to +1 (all videos support misinformation). Positive scores indicate a lean toward misinformation, negative scores indicate a lean toward opposing misinformation, and 0 suggests a set of content that neither supports nor opposes misinformation. A higher score suggests a higher prevalence of misinformation in the SERP.
@@ 7 Results
Here, we compare the prevalence of misinformation in YouTube SERPs across geolocations, topics, and filters. A test of normality revealed that our data is not normal. Thus, we used the non-parametric Mann-Whitney U Test for pairwise comparisons and Kruskal-Wallis tests, followed by the post-hoc Conover-Iman tests with Bonferroni adjustment for multiple comparisons. We provide details of all significance tests in Appendix I. Additionally, we include additional analyses on the misinformation bias scores in search queries and temporal trends of misinformation bias between the US and SA in Appendix J. Lastly, we conducted validation checks to reinforce the robustness of our findings, accounting for the classifier’s 15% error rate and alternative treatments of the COVID-19 origins class in Appendix K.
@@
Among the top-N videos, the top-10 videos in the SERPs have the highest misinformation bias scores. We observed that the top-10 videos in the SERPs have the highest misinformation bias scores for both the US and SA (see Figure 2). This suggests that misinformative content is more prevalent in top-10 search results than lower-ranked results. Overall, the misinformation bias scores across the top-N videos in the SERPs are consistently near -0.05, indicating an general trend towards neutral-to-opposing misinformation in both the US and SA. However, we observe that nearly a third of the top-10 search results were misinformative: 31.55% of the search results supported COVID-19 misinformation, 36.03% were opposing, and 32.42% belonged to the remaining classes.
Bots in SA encountered significantly more misinformative SERPs than bots in the US for the top-10 and top-20 videos. For the top-10 videos, the effect size r was 0.49 (see Appendix Table 8), indicating a “medium effect” (Cohen 1988). This effect size indicates that our observed differences moderately carry practical significance, suggesting that geolocation influences the algorithmic behaviors of YouTube search — in our case, resulting in statistically significantly more misinformative SERPs in SA compared to the US. We discuss the potential implications of the observed differences in §8.
When considering top-40 and top-50 search results, bots in the US encountered significantly more misinformative SERPs than bots in SA. This suggests that misinformative videos are still present in the US, but are downranked to lower positions in SERPs, indicating that YouTube’s content moderation strategy in the US mitigates misinformation through algorithmic adjustments—downranking such content and reducing their visibility rather than outright removal. However, it is important to note that 95% of user traffic is directed towards the first page of search results (Kaye 2013), where the top-10 results reside, making it more likely for users to interact with these results than with lower-ranked results. As a result, the difference in the top-40 and top-50 results may have less impact on user exposure compared to the top-10 results, where SA users are more likely to encounter misinformation. Since video viewers are most likely to engage with the top-10 videos in SERPs, our subsequent analysis focuses on these search results.
@@
Bots in SA encountered significantly more misinformative SERPs than bots in the US for 3 topics. In Figure 3, the topics “5G claims” (r=0.80), “Bill Gates Claims” (r=0.44), and “Vaccine Content Claims” (r=0.59) exhibited statistically significant differences in misinformation bias between the US and SA, indicating medium to large effect sizes. These effect sizes imply practical significance, suggesting that users in SA may encounter significantly more misinformative SERPs than users in the US for these topics. Of particular concern is the presence of these differences in topics related to public health, such as “Vaccine Content Claims.“ In contrast, bots in the US encountered significantly more misinformative SERPs than bots in the SA for a single topic—“Spread of Virus.” We discuss the implications of these findings in §8.
@@
Topics “Biological Weapon” and “Lab Leak Theory” exhibited positive scores across all search filters in both countries (see Figure 4). Given that positive scores indicate that the SERPs lean toward misinformative content, this finding suggests the problematic nature of these topics, with bots exposed to misinformative SERPs regardless of their geolocations or the search filters used.
Topics “5G Claims,” “Spread of Virus,” and “Vaccine Content Claims” displayed negative scores across all search filters in both countries. This suggests that SERPs for these topics generally oppose misinformation in both the US and SA. This result may be attributed to YouTube’s targeted content moderation efforts on these topics (Hern 2020), potentially highlighting some successes in combating misinformation on its platform.
Three topics showed negative scores in SERPs sorted by “Relevance” and “View Count,” but positive scores when sorted by “Upload Date” and “Rating.” We observe this pattern for “Bill Gates Claims, “Treatment of Virus,” and “Population Control” topics, indicating that bots encounter more misinformative SERPs when searching for newly uploaded or highly rated videos for such topics.
@@
In contrast, “Relevance” and “View Count” filters showed negative scores in both countries. This suggests that SERPs sorted by relevant or most-viewed videos lean towards opposing misinformation. This may suggest that YouTube’s content moderation efforts prioritize relevant and highly viewed videos, which are likely to be surfaced by the search engine or have gained viewer attention.
Bots in SA received significantly more misinformative SERPs than bots in the US when sorting the results by “Relevance.” As shown in Figure 5, the effect size of r=0.86 indicates a large effect, suggesting that users in SA may encounter more misinformative SERPs than their US counterparts when using the default “Relevance” filter.
However, bots in the US received significantly more misinformative SERPs than those in SA when sorting the results by “Rating.” The effect size r=0.48 indicates a medium effect. While this effect size is not as pronounced as when sorting by “Relevance” (r=0.86), it suggests that the observed differences have moderate practical significance. It is important to note that YouTube’s default filter is “Relevance,” suggesting that users are more likely to engage with SERPs sorted by this criterion. Consequently, the observed differences with the “Rating” filter may have a lesser impact on users compared to the “Relevance” filter, where bots in SA received significantly more misinformative SERPs than bots in the US.
@@ 8 Discussion
Auditing YouTube’s search engine for COVID-19 misinformation is urgently needed to ensure algorithmic accountability, protect public health, and develop more responsible web. Despite being designated a global pandemic in 2020, COVID-19 remains a public health threat (Colarossi 2024). While YouTube has been a valuable tool for finding health information during the pandemic, it has also disseminated harmful misinformation, impeding public health efforts and vaccine hesitancy (Do Nascimento et al. 2022).
@@
Furthermore, our results indicate that bots in SA were exposed to significantly more misinformative SERPs compared to bots in the US for the topics “5G Claims,” “Bill Gates Claims,” and “Vaccine Content Claims” (Figure 3). Notably, South Africa experienced distressing incidents during the pandemic, such as the burning of 5G towers fueled by a conspiracy theory linking them to the spread of COVID-19 (Reuters 2021). Concurrently, South Africa grappled with many public health challenges, including a vaccine supply shortage and a deep distrust of medical authorities contributing to vaccine hesitancy (Chutel and Fisher 2021). Given these circumstances, the heightened exposure of SA users to misinformation compared to US users on these topics could potentially exacerbate societal issues in SA, including vaccine hesitancy and undermining public health efforts.
Within the US, our analysis revealed that bots in Houston, Texas, received significantly more misinformative SERPs than those in Los Angeles, California. Prior work has shown that conservatives were more susceptible to believing falsehoods and conspiracy beliefs (Garrett and Bond 2021). Since Houston is in a conservative-leaning state (Texas), promoting significantly more misinformative search results to conservative-leaning users could exacerbate harm by reinforcing or fostering beliefs in COVID-19 misinformation.
However, political leaning alone is unlikely to explain the differences observed within the US. Other factors, such as regional search behaviors, public health policies, and online media use policies may also play a role. The observed geographic differences within the US, as well as between the US and SA, highlight important societal and policy implications for platforms like YouTube. These platforms should assess whether current moderation policies sufficiently address geographic disparities in the prevalence of COVID-19 misinformation in search results. To mitigate these disparities, platforms could analyze region-specific search behaviors and trends to adjust rankings or visibility for content likely to contain misinformation. Additionally, regionally tailored counter-misinformation efforts—such as promoting public health campaigns in high-risk areas—could help curb misinformation across diverse regions.
@@
Historically, past audits have successfully generated awareness about these issues, creating pressure on platforms to improve their algorithm (Raji and Buolamwini 2019). Similarly, we hope our work drives meaningful changes on YouTube and inspires future audits in the Global South.
@@ 9 Limitations and Future Directions
Our work is not without limitations. Our audit used only English search queries to maintain consistency in search query language across both countries. Future studies can audit YouTube’s search results by utilizing queries in diverse regional and local languages within the countries. In addition, our study focuses on YouTube’s search engine; future studies can investigate the impacts of geolocation on the recommendation systems on the home page and video page. As the COVID-19 pandemic remains a public health threat globally, our study focuses on COVID-19 misinformation on YouTube. However, other domains, such as climate change misinformation, warrants attention—our methods, including the experimental setup, can be adapted to investigate misinformation beyond COVID-19.
In our study, we employ sock-puppet bots to emulate the actions of real users, which avoids contaminating real users’ search and recommendation results (explained further in §Ethical Statement). However, it is important to note that while we can analyze what happens when a hypothetical user searches on YouTube, we cannot determine the prevalence of misinformation for real users in South Africa compared to the United States. To address this limitation, future work could ethically and carefully conduct crowdsourced audits (Juneja, Bhuiyan, and Mitra 2023) to gather real user data and compare misinformation prevalence in these different geographic regions.
Given the challenges of conducting audits in the Global South, our audit focuses on a country from the Global North and one from the Global South. Future works should conduct audits in multiple countries within the Global South. Our work uses the misinformation bias score (Hussein, Juneja, and Mitra 2020), capturing the amount of misinformation and the video’s rank but not its relevance. Future works can use metrics like the Convex Aggregated Measure (Lioma, Simonsen, and Larsen 2017), which uses relevance and credibility in ranked lists.
Although we showed in §5 that our classifier performs comparably to YouTube models tackling related issues, we acknowledge that our classifier has an error rate of 15%, potentially influencing downstream misinformation analysis. We partly address this limitation by conducting additional analyses, which provided further evidence of the reliability of our findings despite the error rate of our classifier (see subsection “Validation of Results” in Appendix K).